Kyle is 12 years old. His age is 2 more than half the age of Steven. Find the age of Steven.
Answer by Cromlix(4381)

Hi there,
Kyle is 12 years old. His age is 2 more
than half the age of Steven.
Kyle = 12
12 = 2 + 1/2 age of Steven
12 - 2 = 1/2 age of Steven
10 = 1/2 age of Steven
Multiply both sides by 2
20 = 1/2 (2) age of Steven.
Steven's age is 20 years old.
Hope this helps :-)